# Setting up for failure: automatic discovery of the neural mechanisms of cognitive errors

### v1

@@ Appendix C Supplementary results for feature-cued networks
As mentioned in §4, while index-cued networks enjoyed a plane alignment trajectory that resembled the data with just the introduction of swap errors, probe-cued networks required higher swap errors for proximal distractors to do so. Here, we quantify this improvement in matching trends found in the data. We fitted two models to the cosine similarity trajectories data from each network variant: (1) a four-parameter sigmoid function (slope, x-offset, y-offset, y-scale) and (2) a flat line representing the temporal mean (Figure 8). We fit these to all cosine similarity data in Figure 4D, aggregating across location distance (colours of lines). This approach was motivated by the experimental methodology in Panichello2021, where the authors fitted sigmoid functions to the cosine similarity over time. We adopted this as a quantitative benchmark for evaluating whether our networks produce neural dynamics that match the temporal profile observed in biological data. The comparison to a flat line serves as a control. While the constant swap rate network produces temporally stable alignments that are fit well by a sigmoid with near-zero slope, this does not match the characteristic temporal dynamics reported in the experimental results. By quantifying the mean squared error difference between sigmoid and flat line fits, we can assess which networks exhibit the dynamic temporal evolution of representational alignment seen in the biological data. These results are reported in Table 1, in the order of subplots in Figure 4D.
@@ Appendix D Training details
The network trained to produce swap errors are optimised on a joint loss function. The primary terms of this loss function are the stepwise denoising errors (see below; Ho2020), with the transition kernel mean provided by discretising the continuous RNN dynamics (equation 3) projected to the behavioural subspace. Unlike traditional DDPMs, the transition kernel takes in an $n$ dimensional rate vector, and outputs a $2$ dimensional denoised mean in the behavioural subspace. Without projection $W_{x}$, this dictates the full network dynamics throughout the diffusion period (equation 11), but is followed by a teacher forcing step in the diffusion period during training (algorithm 2). As per the advice of seminal work on DDPMs, we weight each term of this loss equally, even though this violates its original formulation as an ELBO Ho2020.
The loss function also penalises the deviation of the distribution of the behavioural subspace activity at cue offset $\bm{x}_{0}$ from the base (unit normal) distribution across many trials. This is a necessary starting point for the denoising process Ho2020 - because the trajectory is not teacher-forced prior to cue offset, we must explicitly regularise it. We achieve this by applying an L2/Frobenius penalty to the deviation of the first two moments of activity at the start of denoising (i.e. the first timestep of delay 2) from this target distribution Soo2022. Finally, we apply an L2 regulariser on neural activity across the trial, including prior to the denoising period Yang2019, Stroud2023.
Normally, training DDPM parameters across different timesteps, i.e. $p_{\phi}(\cdot|\cdot^{\prime},\tau)$ can be done in parallel across $\tau$: data samples $\bm{x}_{T}$ are noised to a valid sample of $q(\bm{x}_{\tau}|\bm{x}_{T})$ with the noising process accumulated over timesteps, then the DDPM objective can be calculated independently of the noised state at other timesteps. This is because DDPMs are Markovian - trained parameters $\phi$ can denoise a sample independently of previous timesteps. This is not possible when training RNNs to perform memory tasks. The necessary computations - memory maintenance during the delay period and retrieval during the cueing period - are implemented via non-linear dynamical motifs [Vyas2020, Driscoll2022, Versteeg2025] which are also driven by the remainder of the neural state space which is not projected to the response space. As such, training RNNs requires simulation of full task trials in sequence. As neural activity trajectories will likely veer out of the sequence of distributions defined by the noising process $q$, this sequential training risks detrimentally biasing the training distribution. To solve this, we use teacher-forcing to ensure the correct distribution of training targets is satisfied (Appendix D; Williams1989).
@@ D.1 Changing optimisation window for task-optimal networks
Note that only the final network rate vector $\bm{r}_{T}$ is behaviourally relevant, as it is used to evaluate the colour estimate of the network for that trial (Appendix A). We refer to the number of timesteps $T^{*}$ before this critical timestep which the optimisation kicks in as the optimisation window. We maintained the same duration and neural noise schedules of each trial period across all networks (Appendix G). For a different, simpler VWM task with a single, variable delay duration, it has previously been shown that adjusting the optimisation window alters the neural coding schemes used for VWM Stroud2023. Specifically, it affected the cross-temporal decodability of task variables in storage for the duration of the delay before the estimate is produced. In the absence of neural evidence to compare to for the neural coding scheme across this delay Panichello2021, we presented the extreme case of $T^{*}=T$ Figure 4E (left), i.e. the MSE loss function in equation 12 was applied to all timesteps of the second delay. Figure 10 shows the equivalent plot for multiple task-optimal networks with different optimisation windows. In all cases, All estimates were tightly bound to the circle in the behavioural subspace, and no swap errors were observed (Figure 3C [top]). Again, we see that the hallmarks of neural data are not captured, although in networks which are afforded time before being penalised for behavioural deviation, there is a shared alignment in the target-distractor colour planes during the optimisation window. Regardless, without swap errors, it is difficult to make connections to behavioural data, as was the original aim of our novel training method. Investigating the link between this property of neural activity and optimal information loading is key to understanding both the mechanism of memory retrieval, and the dynamical source of swap errors during this time (§5), and is left to future works.
@@ Appendix E Neural signatures of swap errors
Figure 11 provides an extension to the alignment results of Figure 5C (left). There, the item locations are kept fixed, as well as one item’s colour. The response-time hidden activity $\bm{m}_{T}$, when projected to its first three principal components (PCs), forms a characteristic 2 ring structure. These PCs are shared across initial item locations, as long as the fixed colour is the same (blue histogram). This is not the case when the colours are dissimilar (red histogram). Figure 11 provides the mean value of this histogram for a wider range of fixed colour pairs. We see a cyclical structure of PC3 alignment with respect to the fixed stimulus colour. This provides preliminary evidence of a toroidal structure, which would have to be verified with further topological data analysis Cueva2021.
@@ F.1 Task and Generative Model
We generated synthetic behavioural data using the Bayesian causal inference model from wei_cue_combination, which has been successfully fit to human psychophysical data in this task. The generative model assumes that visual and auditory sources ($z_{v}$ and $z_{a}$) are drawn independently from a Gaussian prior with standard deviation $\sigma_{p}$. On each trial, these sources either share a common cause ($C=1$) with probability $p_{\text{common}}$, or are independent ($C=0$). When $C=1$, both cues are generated from the same underlying location; when $C=0$, they are generated independently. The cues are corrupted by Gaussian sensory noise with standard deviations $\sigma_{v}$ and $\sigma_{a}$ respectively (Figure 12c). The ideal observer computes the posterior distribution over $s_{a}$ by marginalising over the latent causal structure:

### v2

@@ Appendix C Supplementary results for feature-cued networks
As mentioned in §4, while index-cued networks enjoyed a plane alignment trajectory that resembled the data with just the introduction of swap errors, probe-cued networks required higher swap errors for proximal distractors to do so. Here, we quantify this improvement in matching trends found in the data. We fitted two models to the cosine similarity trajectories data from each network variant: (1) a four-parameter sigmoid function (slope, x-offset, y-offset, y-scale) and (2) a flat line representing the temporal mean (Figure 8). We fit these to all cosine similarity data in Figure 4D, aggregating across location distance (colours of lines). This approach was motivated by the experimental methodology in Panichello and Buschman [2021], where the authors fitted sigmoid functions to the cosine similarity over time. We adopted this as a quantitative benchmark for evaluating whether our networks produce neural dynamics that match the temporal profile observed in biological data. The comparison to a flat line serves as a control. While the constant swap rate network produces temporally stable alignments that are fit well by a sigmoid with near-zero slope, this does not match the characteristic temporal dynamics reported in the experimental results. By quantifying the mean squared error difference between sigmoid and flat line fits, we can assess which networks exhibit the dynamic temporal evolution of representational alignment seen in the biological data. These results are reported in Table 1, in the order of subplots in Figure 4D.
@@ Appendix D Training details
The network trained to produce swap errors are optimised on a joint loss function. The primary terms of this loss function are the stepwise denoising errors (see below; Ho et al. [2020]), with the transition kernel mean provided by discretising the continuous RNN dynamics (equation 3) projected to the behavioural subspace. Unlike traditional DDPMs, the transition kernel takes in an $n$ dimensional rate vector, and outputs a $2$ dimensional denoised mean in the behavioural subspace. Without projection $W_{x}$, this dictates the full network dynamics throughout the diffusion period (equation 11), but is followed by a teacher forcing step in the diffusion period during training (algorithm 2). As per the advice of seminal work on DDPMs, we weight each term of this loss equally, even though this violates its original formulation as an ELBO Ho et al. [2020].
The loss function also penalises the deviation of the distribution of the behavioural subspace activity at cue offset $\bm{x}_{0}$ from the base (unit normal) distribution across many trials. This is a necessary starting point for the denoising process Ho et al. [2020] - because the trajectory is not teacher-forced prior to cue offset, we must explicitly regularise it. We achieve this by applying an L2/Frobenius penalty to the deviation of the first two moments of activity at the start of denoising (i.e. the first timestep of delay 2) from this target distribution Soo and Lengyel [2022]. Finally, we apply an L2 regulariser on neural activity across the trial, including prior to the denoising period Yang et al. [2019], Stroud et al. [2023].
Normally, training DDPM parameters across different timesteps, i.e. $p_{\phi}(\cdot|\cdot^{\prime},\tau)$ can be done in parallel across $\tau$: data samples $\bm{x}_{T}$ are noised to a valid sample of $q(\bm{x}_{\tau}|\bm{x}_{T})$ with the noising process accumulated over timesteps, then the DDPM objective can be calculated independently of the noised state at other timesteps. This is because DDPMs are Markovian - trained parameters $\phi$ can denoise a sample independently of previous timesteps. This is not possible when training RNNs to perform memory tasks. The necessary computations - memory maintenance during the delay period and retrieval during the cueing period - are implemented via non-linear dynamical motifs [Vyas et al., 2020, Driscoll et al., 2022, Versteeg et al., 2025] which are also driven by the remainder of the neural state space which is not projected to the response space. As such, training RNNs requires simulation of full task trials in sequence. As neural activity trajectories will likely veer out of the sequence of distributions defined by the noising process $q$, this sequential training risks detrimentally biasing the training distribution. To solve this, we use teacher-forcing to ensure the correct distribution of training targets is satisfied (Appendix D; Williams and Zipser [1989]).
@@ D.1 Changing optimisation window for task-optimal networks
Note that only the final network rate vector $\bm{r}_{T}$ is behaviourally relevant, as it is used to evaluate the colour estimate of the network for that trial (Appendix A). We refer to the number of timesteps $T^{*}$ before this critical timestep which the optimisation kicks in as the optimisation window. We maintained the same duration and neural noise schedules of each trial period across all networks (Appendix G). For a different, simpler VWM task with a single, variable delay duration, it has previously been shown that adjusting the optimisation window alters the neural coding schemes used for VWM Stroud et al. [2023]. Specifically, it affected the cross-temporal decodability of task variables in storage for the duration of the delay before the estimate is produced. In the absence of neural evidence to compare to for the neural coding scheme across this delay Panichello and Buschman [2021], we presented the extreme case of $T^{*}=T$ Figure 4E (left), i.e. the MSE loss function in equation 12 was applied to all timesteps of the second delay. Figure 10 shows the equivalent plot for multiple task-optimal networks with different optimisation windows. In all cases, All estimates were tightly bound to the circle in the behavioural subspace, and no swap errors were observed (Figure 3C [top]). Again, we see that the hallmarks of neural data are not captured, although in networks which are afforded time before being penalised for behavioural deviation, there is a shared alignment in the target-distractor colour planes during the optimisation window. Regardless, without swap errors, it is difficult to make connections to behavioural data, as was the original aim of our novel training method. Investigating the link between this property of neural activity and optimal information loading is key to understanding both the mechanism of memory retrieval, and the dynamical source of swap errors during this time (§5), and is left to future works.
@@ Appendix E Neural signatures of swap errors
Figure 11 provides an extension to the alignment results of Figure 5C (left). There, the item locations are kept fixed, as well as one item’s colour. The response-time hidden activity $\bm{m}_{T}$, when projected to its first three principal components (PCs), forms a characteristic 2 ring structure. These PCs are shared across initial item locations, as long as the fixed colour is the same (blue histogram). This is not the case when the colours are dissimilar (red histogram). Figure 11 provides the mean value of this histogram for a wider range of fixed colour pairs. We see a cyclical structure of PC3 alignment with respect to the fixed stimulus colour. This provides preliminary evidence of a toroidal structure, which would have to be verified with further topological data analysis Cueva et al. [2021].
@@ F.1 Task and Generative Model
We generated synthetic behavioural data using the Bayesian causal inference model from Beierholm et al. [2007], which has been successfully fit to human psychophysical data in this task. The generative model assumes that visual and auditory sources ($z_{v}$ and $z_{a}$) are drawn independently from a Gaussian prior with standard deviation $\sigma_{p}$. On each trial, these sources either share a common cause ($C=1$) with probability $p_{\text{common}}$, or are independent ($C=0$). When $C=1$, both cues are generated from the same underlying location; when $C=0$, they are generated independently. The cues are corrupted by Gaussian sensory noise with standard deviations $\sigma_{v}$ and $\sigma_{a}$ respectively (Figure 12c). The ideal observer computes the posterior distribution over $s_{a}$ by marginalising over the latent causal structure: